# IFT-UAM/CSIC-25-108

### v1

@@ I Variation in chi-squared weights and model-selection criteria
$\mathrm{Tay2}][a^{2}][-]$ 0.82 0.0384 0.1448(6)
$\mathrm{Tay2}][a^{2}][\beta>3.40]$ 0.78 0.0162 0.1446(8)
$\mathrm{Tay2}][a^{2}][\beta>3.46]$ 0.73 0.0028 0.1445(10)
$\mathrm{Tay2}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.77 0.0065 0.1447(6)
$\mathrm{Tay2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.65 0.0009 0.1444(10)
$\mathrm{Tay2}][a^{2}][m_{\pi}<350\,\mathrm{MeV}]$ 0.52 0.0001 0.1445(7)
$\mathrm{Tay2}][a^{2}][m_{\pi}L>4.1]$ 0.81 0.0028 0.1449(7)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.83 0.0390 0.1449(6)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.77 0.0152 0.1447(8)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.72 0.0029 0.1446(10)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\,\mathrm{MeV}]$ 0.77 0.0068 0.1447(6)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.65 0.0009 0.1444(11)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\,\mathrm{MeV}]$ 0.52 0.0001 0.1446(8)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.81 0.0027 0.1450(7)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.88 0.0533 0.1444(7)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.80 0.0142 0.1443(9)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.71 0.0016 0.1443(10)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.81 0.0054 0.1444(7)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.60 0.0004 0.1442(11)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<350\,\mathrm{MeV}]$ 0.44 0.0000 0.1443(8)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.89 0.0044 0.1444(8)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][-]$ 0.84 0.0246 0.1450(6)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.78 0.0100 0.1448(9)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.70 0.0016 0.1447(10)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.78 0.0049 0.1449(6)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.61 0.0004 0.1446(11)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<350\,\mathrm{MeV}]$ 0.42 0.0000 0.1446(8)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.83 0.0020 0.1452(7)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][-]$ 0.86 0.0371 0.1441(9)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.79 0.0077 0.1438(14)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.65 0.0008 0.1444(18)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.77 0.0030 0.1442(12)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.89 0.0024 0.1438(14)
@@
$\mathrm{Tay4}][a^{2}][-]$ 0.79 0.0179 0.1446(8)
$\mathrm{Tay4}][a^{2}][\beta>3.40]$ 0.73 0.0053 0.1445(10)
$\mathrm{Tay4}][a^{2}][\beta>3.46]$ 0.67 0.0011 0.1444(11)
$\mathrm{Tay4}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.73 0.0030 0.1446(8)
$\mathrm{Tay4}][a^{2}][m_{\pi}L>4.1]$ 0.79 0.0015 0.1455(12)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.85 0.0185 0.1443(8)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.76 0.0054 0.1442(11)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.62 0.0007 0.1442(12)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.77 0.0028 0.1443(8)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.90 0.0028 0.1450(12)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][-]$ 0.80 0.0132 0.1447(8)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.73 0.0036 0.1446(10)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.64 0.0006 0.1446(12)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.75 0.0025 0.1447(8)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.79 0.0011 0.1456(12)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][-]$ 0.82 0.0134 0.1441(9)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.72 0.0029 0.1438(14)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.58 0.0003 0.1445(17)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.72 0.0014 0.1442(12)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.97 0.0057 0.1436(15)
@@
$\chi\mathrm{SU(3)}][a^{2}][-]$ 0.13 0.0051 0.1455(7)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40]$ 0.24 0.0235 0.1444(10)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.46]$ 0.20 0.0112 0.1437(12)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.15 0.0007 0.1453(8)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.11 0.0001 0.1455(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.12 0.0070 0.1455(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.23 0.0231 0.1444(10)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.20 0.0111 0.1436(12)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\,\mathrm{MeV}]$ 0.16 0.0020 0.1453(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.12 0.0002 0.1455(8)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.22 0.0656 0.1447(8)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.39 0.1906 0.1437(11)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.45 0.1062 0.1428(12)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.21 0.0033 0.1446(9)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.27 0.0042 0.1445(9)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][[-]$ 0.17 0.0078 0.1459(7)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.24 0.0250 0.1449(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.20 0.0024 0.1443(12)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.20 0.0031 0.1457(8)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.16 0.0005 0.1461(8)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.19 0.0317 0.1450(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.35 0.0753 0.1435(13)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.45 0.0309 0.1420(16)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.21 0.0020 0.1450(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.23 0.0012 0.1443(12)
@@
$\chi\mathrm{SU(2)}][a^{2}][-]$ 0.15 0.0052 0.1457(7)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40]$ 0.24 0.0086 0.1445(10)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.46]$ 0.19 0.0012 0.1439(12)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.11 0.0001 0.1453(9)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.12 0.0001 0.1459(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.15 0.0057 0.1457(8)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.22 0.0085 0.1445(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.18 0.0012 0.1439(12)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\,\mathrm{MeV}]$ 0.12 0.0001 0.1453(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.12 0.0001 0.1459(9)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.26 0.0298 0.1449(8)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.38 0.1536 0.1439(11)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.43 0.0299 0.1431(12)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.17 0.0006 0.1446(9)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.29 0.0026 0.1449(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][[-]$ 0.22 0.0159 0.1461(7)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.25 0.0201 0.1450(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.20 0.0018 0.1446(12)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.16 0.0005 0.1456(9)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.16 0.0003 0.1464(9)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.24 0.0143 0.1452(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.36 0.0484 0.1436(13)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.42 0.0158 0.1423(17)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.16 0.0007 0.1449(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.27 0.0009 0.1446(13)

### v2

@@ I Variation in chi-squared weights and model-selection criteria
$\mathrm{Tay2}][a^{2}][-]$ 0.74 0.0380 0.1445(6)
$\mathrm{Tay2}][a^{2}][\beta>3.40]$ 0.71 0.011 0.1443(7)
$\mathrm{Tay2}][a^{2}][\beta>3.46]$ 0.70 0.0018 0.1440(9)
$\mathrm{Tay2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.76 0.0097 0.1443(6)
$\mathrm{Tay2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.67 0.0009 0.1439(9)
$\mathrm{Tay2}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.57 0.0001 0.1444(7)
$\mathrm{Tay2}][a^{2}][m_{\pi}L>4.1]$ 0.70 0.0016 0.1445(6)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.72 0.0291 0.1445(6)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.68 0.0090 0.1443(7)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.69 0.0024 0.1440(9)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.75 0.0079 0.1443(6)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.68 0.0008 0.1439(9)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.57 0.0001 0.1444(7)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.68 0.0013 0.1445(7)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.86 0.0531 0.1441(6)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.81 0.0129 0.1439(8)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.72 0.0016 0.1438(9)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.81 0.0079 0.1441(6)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.62 0.0005 0.1437(9)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<350\;\rm{MeV}]$ 0.52 0.0000 0.1442(7)
$\mathrm{Tay2}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.85 0.0027 0.1440(7)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][-]$ 0.79 0.0268 0.1448(6)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.75 0.0080 0.1446(8)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.71 0.0014 0.1443(9)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.78 0.0075 0.1445(6)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.63 0.0004 0.1441(10)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.50 0.0000 0.1445(7)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.77 0.0011 0.1449(7)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][-]$ 0.87 0.0413 0.1437(8)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.78 0.0084 0.1434(12)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.67 0.0008 0.1439(15)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.77 0.0039 0.1438(11)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.84 0.0016 0.1435(12)
@@
$\mathrm{Tay4}][a^{2}][-]$ 0.72 0.0157 0.1442(8)
$\mathrm{Tay4}][a^{2}][\beta>3.40]$ 0.68 0.0047 0.1440(9)
$\mathrm{Tay4}][a^{2}][\beta>3.46]$ 0.66 0.0008 0.1438(10)
$\mathrm{Tay4}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.72 0.0041 0.1440(8)
$\mathrm{Tay4}][a^{2}][m_{\pi}L>4.1]$ 0.68 0.0007 0.1450(10)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.84 0.0199 0.1439(8)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.76 0.0054 0.1437(9)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.66 0.0006 0.1436(10)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.77 0.0026 0.1439(8)
$\mathrm{Tay4}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.87 0.0012 0.1447(10)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][-]$ 0.79 0.0159 0.1444(8)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.72 0.0037 0.1442(9)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.68 0.0006 0.1440(10)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.75 0.0023 0.1442(8)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.73 0.0005 0.1452(10)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][-]$ 0.83 0.0116 0.1437(9)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.73 0.0032 0.1434(12)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.62 0.0003 0.1440(15)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.71 0.0017 0.1439(11)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.94 0.0021 0.1434(13)
@@
$\chi{\rm SU(3)}][a^{2}][-]$ 0.10 0.0054 0.1455(7)
$\chi{\rm SU(3)}][a^{2}][\beta>3.40]$ 0.23 0.0237 0.1445(9)
$\chi{\rm SU(3)}][a^{2}][\beta>3.46]$ 0.19 0.0050 0.1437(11)
$\chi{\rm SU(3)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.14 0.0016 0.1454(7)
$\chi{\rm SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.11 0.0002 0.1456(7)
$\chi{\rm SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.11 0.0061 0.1456(7)
$\chi{\rm SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.21 0.0231 0.1445(9)
$\chi{\rm SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.19 0.0049 0.1437(11)
$\chi{\rm SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.15 0.0021 0.1454(7)
$\chi{\rm SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.11 0.0002 0.1456(7)
$\chi{\rm SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.19 0.0625 0.1447(7)
$\chi{\rm SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.35 0.3014 0.1438(9)
$\chi{\rm SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.40 0.0481 0.1429(11)
$\chi{\rm SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.19 0.0041 0.1446(8)
$\chi{\rm SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.26 0.0028 0.1446(8)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][[-]$ 0.17 0.0288 0.1460(7)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.23 0.0401 0.1449(9)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.21 0.0063 0.1443(10)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.19 0.0013 0.1458(7)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.16 0.0011 0.1462(7)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.18 0.0291 0.1451(9)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.32 0.0745 0.1436(11)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.41 0.0614 0.1421(13)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.19 0.0085 0.1451(9)
$\chi{\rm SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.22 0.0025 0.1445(10)
@@
$\chi{\rm SU(2)}][a^{2}][-]$ 0.12 0.0047 0.1457(7)
$\chi{\rm SU(2)}][a^{2}][\beta>3.40]$ 0.20 0.0096 0.1446(9)
$\chi{\rm SU(2)}][a^{2}][\beta>3.46]$ 0.18 0.0015 0.1439(10)
$\chi{\rm SU(2)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.10 0.0002 0.1453(8)
$\chi{\rm SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.11 0.0001 0.1459(8)
$\chi{\rm SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.13 0.0054 0.1458(7)
$\chi{\rm SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.21 0.0087 0.1446(9)
$\chi{\rm SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.17 0.0014 0.1439(10)
$\chi{\rm SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.10 0.0002 0.1453(8)
$\chi{\rm SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.12 0.0001 0.1460(8)
$\chi{\rm SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.21 0.0429 0.1449(8)
$\chi{\rm SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.34 0.0574 0.1439(10)
$\chi{\rm SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.36 0.0156 0.1431(11)
$\chi{\rm SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.15 0.0006 0.1447(9)
$\chi{\rm SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.26 0.0023 0.1449(8)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][[-]$ 0.19 0.0180 0.1461(7)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.22 0.0070 0.1451(9)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.19 0.0014 0.1446(10)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.14 0.0004 0.1457(8)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.15 0.0004 0.1464(8)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.20 0.0121 0.1452(9)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.32 0.0326 0.1437(11)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.36 0.0315 0.1424(13)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.14 0.0004 0.1450(9)
$\chi{\rm SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.22 0.0008 0.1447(11)